### v1

# Supervised Learning-enhanced Multi-Group Actor Critic for Live-stream Recommendation

Abstract.
Reinforcement Learning (RL) has been widely applied in recommendation systems to capture users’ long-term engagement, thereby improving dwelling time and enhancing user retention. In the context of a short video & live-stream mixed recommendation scenario, the live-stream recommendation system (RS) decides whether to inject at most one live-stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live-stream injection policy for accurate live-stream allocation. However, traditional RL algorithms often face divergence and instability problems, and these issues are even more pronounced in our scenario. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods but also exhibits enhanced stability in online recommendation scenarios.
Reinforcement Learning, Recommendation System, Variance Reduction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
Figure 1. Structure of a short video & live-stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live-stream RS.
1. Introduction
Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards across various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017). In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023). Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
We consider a challenging sequential decision-making scenario in a short video & live-stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live-stream recommendation system, a short video recommendation system, and a blending server. For each user request at timestamp $t$, the live-stream recommendation system decides whether to inject the recommended live-stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live-stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live-stream injection policy that maximizes users’ long-term engagement with live-streams, which can be modeled as an infinite request-level Markov Decision Process (MDP) (Puterman, 1990).
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning (RL) algorithms (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario. A primary possible reason for this is the drastic fluctuations in both the live-stream supply scale and user interaction behaviors over time. As shown in Fig. 2a, the curves of live-stream room count and live-stream viewer count vary significantly on a daily basis, making it difficult for the RL algorithm to learn optimal policies. Furthermore, as illustrated in Fig. 2b, the distributions of user interaction labels (e.g., live-stream and short video watch time) are highly noisy and exhibit large variance. RL models often struggle to learn effective policies from data that exhibit such high variance across both time and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
Figure 2. Data distributions of online live-stream room (or viewer) count, live-stream watch time and short video watch time.
To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC. Given the significant variation in users’ interest in live-stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks. Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation. Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
The main contributions of this paper are as follows:
  • We introduce SL-MGAC, a novel RL model designed to reduce prediction variance while enhancing the stability of both offline training and online inference.
  • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q-value normalization.
  • We integrate multi-task supervised learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
  • We successfully deploy the SL-MGAC model in a challenging short video and live-stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
2. Problem Formulation
As shown in Fig. 1, in the short video & live-stream mixed recommendation system, the live-stream recommendation system is treated as an agent that interacts with diverse users and receives user feedback over time. The optimal live-stream injection control problem is modeled as an infinite request-level MDP to maximize the cumulative reward. Formally, we define the live-stream injection MDP as a tuple of five elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right)$:
  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live-stream watch history, short video watch history), and item features (e.g., live-stream ID, author ID, author gender, online viewer count). We limit the length of users’ live-stream and short video history lists by timestamp order, keeping only the top 50 items.
  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live-stream in response to a user’s request. We define action $a$ as a binary variable, where $a=1$ means injecting a live-stream, and $a=0$ means not injecting it.
  • Transition Probability $\mathcal{P}$: The transition probability is denoted as $p(s_{t+1}|s_{t},a_{t})$, determined by the environment.
  • Reward Function $\mathcal{R}$: The reward function $\mathcal{R}$ is a mapping from state $s_{t}$ and action $a_{t}$ at timestamp $t$, which can be formulated as $r(s_{t},a_{t}):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$.
  • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the computation of cumulative reward. Typically, we set $\gamma<1$ during model training.
Figure 3. Overall framework of the SL-MGAC algorithm. The SL (RL) MG-SD Module is short for the Multi-Group State Decomposition Module for supervised reward learning and critic learning.
Overall, the optimization objective of aforementioned optimal live-stream injection control problem is shown as follows:
(1) $\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
where $\pi(a_{t}|s_{t})$ is the live-stream injection policy to be optimized.
2.1. Reward Function Design
In our mixed short video & live-stream recommendation scenario, we design a reward function based on the time gain of live-stream watch time relative to the average short video watch time in a request:
(2) $r(s_{t},a_{t})=\begin{cases}y_{l}-\frac{1}{B}y_{v},\quad&if\;a_{t}=1\\ \beta&otherwise\\ \end{cases}$
where $y_{l}$ is the live-stream watch time, $y_{v}$ is the total video watch time, $B$ is the number of videos in a request, and $\beta$ is the hyper-parameter in Eq. 2. Note that the recommended list in a request always contains $B$ videos and at most one live-stream. Furthermore, the reward defined above is similar to that of the RCPO (Tessler et al., 2018) algorithm. From the perspective of constrained reinforcement learning, the term $\frac{1}{B}y_{v}$ can be viewed as an implicit cost constraint. In practice, the hyper-parameter $\beta$ is introduced to carefully control the final live-stream injection ratio.
Since the distribution of live-stream (short video) watch time changes drastically over time, as shown in Fig. 2b, it is challenging for an RL agent to optimize the cumulative reward in Eq. 2. To address this, we propose a reward normalization technique as follows:
(3) $r(s_{t},a_{t})=\begin{cases}sigmoid\left(y_{l}-\frac{1}{B}y_{v}\right),\quad&if\;a_{t}=1\\ \beta\in[0,1]&otherwise\\ \end{cases}$
which helps stabilize the network gradient, thereby reducing the risk of model divergence.
3. Proposed Framework
To address the intrinsic divergence and instability issues of reinforcement learning and successfully deploy the RL agent in our mixed short video & live-stream recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC), as shown in Fig. 3. SL-MGAC incorporates a supervised learning-enhanced actor-critic model with a shared user & live-stream feature extraction module, along with independent Multi-Group State Decomposition (MG-SD) modules. For clarity, only one critic network is shown in Fig. 3, while the remaining three critic networks, which use the same architecture, are omitted.
3.1. User & Live-stream Feature Extraction Module
The user & live-stream feature extraction module aims to generate non-linear embedding representations for the state $s_{t}$. First, we use a unified embedding layer to map the user’s static features, live-stream features, historical live-stream list features and short video list features into low-dimensional dense representations. We denote $\mathbf{v}_{U}$, $\mathbf{v}_{L}$, $\left\{\mathbf{e}_{1},\cdots,\mathbf{e}_{M_{l}}\right\}$, $\left\{\mathbf{e}^{\prime}_{1},\cdots,\mathbf{e}^{\prime}_{M_{v}}\right\}$ as the corresponding embedding vectors or sets of embedding vectors, respectively. We then define $\mathbf{h}_{a}=[\mathbf{v}_{U},\mathbf{v}_{L}]$ as the concatenation of $\mathbf{v}_{U}$ and $\mathbf{v}_{L}$.
To aggregate historical live-stream (short video) representations, we introduce the target attention mechanism (Vaswani, 2017; Zhou et al., 2018), which is defined as follows:
(4) $\mathbf{h}_{live}$ $=\sum_{i=0}^{M_{l}}f_{l}(\mathbf{h}_{a},\mathbf{e}_{i})\mathbf{e}_{i}$
$\mathbf{h}_{video}$ $=\sum_{j=0}^{M_{v}}f_{v}(\mathbf{h}_{a},\mathbf{e}^{\prime}_{j})\mathbf{e}_{j}^{\prime}$
where $f_{l}$, $f_{v}$ are different target attention functions, such as a feed-forward network whose output is a normalized score.
After aggregating the historical live-stream and short video features via two separate attention networks, we concatenate all the above embedding vectors to form $\mathbf{h}_{s}=[\mathbf{v}_{U},\mathbf{v}_{L},\mathbf{h}_{live},\mathbf{h}_{video}]$. We then use a shared multi-layer perceptron (MLP) for both the actor and critic networks to obtain the latent representation of state $s_{t}$, i.e. $\mathbf{h}^{\prime}_{s}=f_{MLP}(\mathbf{h}_{s})$. To ensure training stability, we stop the gradient flow through $\mathbf{h}^{\prime}_{s}$ for the actor network and only use the more complex critic networks to optimize $\mathbf{h}^{\prime}_{s}$, as we find that sharing the feature extraction module and the $f_{MLP}$ network causes interference between the policy and critic gradients.
3.2. Multi-Group State Decomposition Module
Since live-stream content is not equally appealing to all users on the short video app Kwai, it is crucial to inject live-streams selectively into the short video feed. Otherwise, excessive live-stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration. In practice, user interaction data for live-streams is sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live-stream injection policy for each individual user.
Therefore, prior information about diverse user groups is essential for enhancing the decision-making accuracy of RL models. A natural approach to distinguish between different user behaviors is to partition users into distinct groups based on their historical activity. Specifically, we categorize users into $K$ disjoint groups according to their live-stream activity level, which is determined by the cumulative live-stream watch time over the past 3 weeks. Users with higher historical watch time are assigned to groups with higher activity levels.
Formally, we present the following theorem, which demonstrates that the proposed multi-group state decomposition can be viewed as a type of variance reduction technique for real-world RL applications.
Theorem 3.1.
Let state space $\mathcal{S}=\bigcup_{i=1}^{K}\mathcal{S}_{i},\forall i\neq j\;\mathcal{S}_{i}\cap\mathcal{S}_{j}=\varnothing$, state $s\in\mathcal{S}$; $r$ be the real reward of state $s$ and action $a$, $\hat{r}$ be the predicted value of $r$ over $\mathcal{S}$, $\hat{r}_{i}$ be the predicted reward of $r$ over $\mathcal{S}_{i}$; then a neural network $f$ with MG-SD module has a lower prediction variance than the neural network $f^{\prime}$ without MG-SD module.
Please refer to Appendix. A for detailed proof of Theorem 3.1. We demonstrate that the MG-SD module eliminates between-group variance, thereby reducing the total variance, which implies that the multi-group state decomposition module contributes to a more stable RL model. Empirically, we also show the effectiveness of this module in enhancing RL model stability in subsequent experiments. We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
3.3. Supervised Learning-enhanced Actor Critic
To alleviate the impact of drastic changes in data distribution that could lead to the divergence or instability of RL models, we propose a supervised learning-enhanced actor critic framework to prevent critic networks from getting trapped in model collapse due to large cumulative errors in the critic learning process.
3.3.1. Layer Normalization
A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern. Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse. Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate the divergence and instability problems.
From the proof in the Appendix D of (Yue et al., 2024), we know that for any input $\mathbf{x}$ and any direction $\mathbf{v}$, if a network applies Layer Normalization to the input, then we have $k_{\rm NTK}(\mathbf{x},\mathbf{x}+\lambda\mathbf{v})\leq C$, where $\lambda>0$, $C$ is a constant and $k_{\rm NTK}$ is the Neural Tangent Kernel (NTK) (Jacot et al., 2018). This theoretically indicates that a network with Layer Normalization is less sensitive to input variations and can maintain stable gradients despite perturbations during model training. Hence, we apply Layer Normalization to the inputs of both the actor and critic networks, as shown in Fig. 3.
3.3.2. Critic Network
Let $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1},r_{t+1})\in\mathcal{D}$ be a training sample from our real-time dataset $\mathcal{D}$. To address the maximization bias problem (Van Hasselt et al., 2016), we employ four critic networks: two current Q-networks $Q_{\phi_{1}}$, $Q_{\phi_{2}}$ and two corresponding target Q-networks $Q^{\prime}_{\phi_{1}}$, $Q^{\prime}_{\phi_{2}}$ in Clipped Double Q-Learning (Fujimoto et al., 2018). The critic learning objective is as follows:
(5) $\mathcal{L}_{Critic}=\sum_{i=1}^{2}\mathbb{E}_{(s,a)\in\mathcal{D}}$ $\left[Q_{\phi_{i}}(s_{t},a_{t})-Q_{label}(s_{t+1})\right]^{2}$
$Q_{label}(s_{t+1})$ $=r(s_{t},a_{t})+\gamma\max_{a_{t+1}}\hat{Q}^{\prime}(s_{t+1},a_{t+1})$
where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use Huber Loss (Hastie et al., 2009) to optimize the above objective.
In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of the critic networks. Therefore, we seamlessly introduce supervised learning (e.g. multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure. Specifically, we split the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
(6) $Q_{\phi_{i}}(s_{t},a_{t})$ $=R_{\theta_{i}}(s_{t},a_{t})+\gamma T_{\xi_{i}}(s_{t},a_{t}),\quad i=1,2$
$Q^{\prime}_{\phi_{i}}(s_{t+1},a_{t+1})$ $=R^{\prime}_{\theta_{i}}(s_{t+1},a_{t+1})+\gamma T^{\prime}_{\xi_{i}}(s_{t+1},a_{t+1}),\quad i=1,2$
where $R_{\theta_{i}}$ and $T_{\xi_{i}}$ are RPN and QRN, $R^{\prime}_{\theta_{i}}$ and $T^{\prime}_{\xi_{i}}$ are target RPN and QRN, respectively. We use separate MLPs to model the RPN and QRN.
Since the distribution of live-stream (short video) watch time changes drastically over time, the time gain reward $r$ in Eq. 2 becomes difficult to learn. In this work, we propose a distribution discretization method to improve reward learning.
Formally, we divide the live-stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live-stream (short video) watch time. Let $y$ denote the actual live-stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$. The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$. Then, we have a linear reconstruction layer to reconstruct $y$ from $\delta$:
(7) $y$ $=y_{st}+\delta\cdot(y_{end}-y_{st})$
which resembles the structure of linear reward shifting (Sun et al., 2022).
Let $o(t)_{l}\in\mathbb{R}^{N_{l}+1}$, $o(t)_{v}\in\mathbb{R}^{N_{v}+1}$ be one-hot vectors representing the time bins in which the real live-stream (short video) watch time of sample $(s_{t},a_{t},r_{t})$ falls. Note that a separate bin is set for the case $a_{t}=0$. Then, the RPN $R_{\theta_{i}}$ can be modeled by multi-task neural networks:
(8) $R_{\theta_{i}}(s_{t},a_{t})$ $=sigmoid\left(F_{\Gamma_{i}}(s_{t},a_{t})-G_{\Theta_{i}}(s_{t},a_{t})\right)$
$F_{\Gamma_{i}}(s_{t},a_{t})$ $={o(t)_{l}}^{T}\left(y_{l}^{st}+f_{\Gamma_{i}}(s_{t},a_{t})\cdot\left(y_{l}^{end}-y_{l}^{st}\right)\right)$
$G_{\Theta_{i}}(s_{t},a_{t})$ $={o(t)_{v}}^{T}\left(y_{v}^{st}+g_{\Theta_{i}}(s_{t},a_{t})\cdot\left(y_{v}^{end}-y_{v}^{st}\right)\right)$
where $y_{l}^{st},y_{l}^{end}\in\mathbb{R}^{N_{l}+1}$ predefined left (right) boundary vectors for the live-stream watch time bin, $y_{v}^{st},y_{v}^{end}\in\mathbb{R}^{N_{v}+1}$ are predefined left (right) boundary vectors for the short video watch time bin. Note that introducing the posterior one-hot vectors $o(t)_{l}$ and $o(t)_{v}$ in Eq. 8 will not affect the calculation of $Q^{\prime}_{\phi_{i}}(s_{t+1},\cdot)$, because we have already reserved $r_{t+1}$ in the dataset $\mathcal{D}$. Therefore, $o(t+1)_{l}$ and $o(t+1)_{v}$ can be easily obtained from $r_{t+1}$ to compute $R^{\prime}_{\theta_{i}}(s_{t+1},\cdot)$.
As shown in Eq. 8, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin. We then reconstruct the predicted live-stream and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively. Finally, we obtain the predicted normalized reward, $R_{\theta_{i}}(s_{t},a_{t})$. To optimize the predicted watch time proportions, we use Huber Loss as follows:
(9) $\mathcal{L}_{SL}=\sum_{i=1}^{2}Huber\_Loss(\delta_{l},F^{\prime}_{\Gamma_{i}})+Huber\_Loss(\delta_{v},G^{\prime}_{\Theta_{i}})$
where $\delta_{l}$, $\delta_{v}$ are the actual proportion labels for live-stream and short video watch times, respectively.
The reason for using Huber loss on the time proportion $\delta$ instead of the actual watch time $y$ is that the time proportion lies within the range $[0,1]$, which results in smaller gradients in the neural network. Additionally, the variance of $\delta$ is much smaller than that of the original watch time $y$. Consequently, the outputs of the learned reward networks ($f_{\Gamma_{i}}$ and $g_{\Theta_{i}}$) will exhibit smaller variances. Hence, reward learning with the discretization of watch time distributions can be viewed as a novel variance reduction technique.
3.3.3. Actor Network
We also incorporate the multi-group state decomposition module into the actor network. The loss function of the actor network is shown below:
(10) $\mathcal{L}_{Actor}$ $=\mathbb{E}_{(s,a)\in\mathcal{D}}\left[-\hat{Q}(s_{t},a_{t})\log p(s_{t},a_{t})\right]$
where $\hat{Q}(s_{t},a_{t})=\min_{i=1,2}Q_{\phi_{i}}(s_{t},a_{t})$, and $p(s_{t},a_{t})$ is the action probability output of the actor network.
Moreover, we observe that large values of $\hat{Q}(s_{t},a_{t})$ may cause instability in actor training, which can ultimately lead to policy deterioration, where $\forall t,\;\pi(a_{t}=1|s_{t})>\pi(a_{t}=0|s_{t})$ or $\pi(a_{t}=1|s_{t})<\pi(a_{t}=0|s_{t})$. To address this, we apply softmax normalization to $\hat{Q}(s_{t},a_{t})$ to obtain $\hat{Q}_{norm}=softmax(\hat{Q}(s_{t},a_{t}))$, and propose a modified actor loss function:
(11) $\mathcal{L}_{Actor}$ $=\mathbb{E}_{(s,a)\in\mathcal{D}}\left[-\hat{Q}_{norm}(s_{t},a_{t})\log p(s_{t},a_{t})\right]$
A similar loss function can be found in AWAC (Nair et al., 2020), and its theoretical results demonstrate that the above loss is equivalent to an implicitly constrained RL problem:
(12) $\pi_{k+1}$ $=\arg\max_{\pi\in\Pi}\mathbb{E}_{\mathbf{a}\sim\pi(\cdot|\mathbf{s})}\left[\hat{Q}^{\pi_{k}}(s_{t},a_{t})\right]$
${\rm s.t.}$ $D_{\rm KL}\left(\pi(\cdot|\mathbf{s})\|\pi_{o}(\cdot|\mathbf{s})\right)\leq\epsilon$
where $D_{\rm KL}$ is the Kullback-Leibler (KL) divergence, and $\pi_{o}$ is the behavior policy derived from the dataset $\mathcal{D}$.
We note that the actor loss in Eq. LABEL:eq:15 is a standard cross-entropy loss function. From the perspective of knowledge distillation (Hinton, 2015), the actor distills policy knowledge from more complex critic networks. The teacher critic networks guide the more lightweight student actor network to adjust and converge to an optimal policy. In practice, we only need to deploy the actor network in the online live-stream recommendation system, which significantly reduces computational complexity and improves real-time respond speed.
Overall, the final loss function of our proposed SL-MGAC algorithm is as follows:
(13) $\mathcal{L}=\mathcal{L}_{Actor}+\mathcal{L}_{Critic}+\mathcal{L}_{SL}$
3.4. Online Exploration and Deployment
For online exploration, we adopt the commonly used $\epsilon$-greedy (Watkins, 1989) strategy:
(14) $\pi_{online}(s_{t})=\begin{cases}{\rm random\;action\;from\;}\mathcal{A}(s_{t}),\quad&if\;\psi<\epsilon\\ \arg\max_{a\in\mathcal{A}(s_{t})}\pi(a_{t}|s_{t}),&{\rm otherwise}\end{cases}$
where $\psi$ is a random number, and $\epsilon$ is maximal exploration probability.
We deploy the SL-MGAC algorithm in our recommendation system, and the overall system architecture is shown in Fig. 4. The online RL agent collects real-time user interaction logs, while the offline model trainer optimizes the SL-MGAC model in an off-policy manner using streaming training data. Moreover, the offline trainer sends the latest model parameters to update the online deployed actor network in real-time.
Figure 4. System Architecture of the SL-MGAC algorithm.
4. Experiments
We conduct both offline evaluation on a real-world dataset collected from our recommendation system and online A/B test experiments with SL-MGAC and the baselines.
4.1. Dataset
Due to the lack of publicly available datasets for recommendation decisions involving live-streaming injections in short video feeds, we collect a dataset from our recommendation system with 10,000 users through random sampling over a full day, ensuring that the proportions of different user groups are similar to those in real online data. In total, we have over 180,000 samples, which are split into training and test sets with a 4:1 ratio.
4.2. Compared Methods
4.2.1. Baselines
We compare our approach with existing non-reinforcement learning and reinforcement learning methods.
  • Learning to Rank (L2R) (Chen and Zhou, 2020). A supervised learning framework that predicts the reward for each action. The action with the maximum reward is selected as the agent’s action.
  • DQN (Mnih et al., 2015). A deep neural network algorithm for Q-learning that introduces the technique of target network updates.
  • BCQ (Fujimoto et al., 2019). A widely used offline reinforcement learning algorithm that adds action restrictions and encourages the agent to behave similarly to the behavior policy.
  • SAC (Haarnoja et al., 2018; Christodoulou, [n. d.]). A classic off-policy reinforcement learning method that maximizes the trade-off between cumulative reward and policy entropy. We use the discrete version of SAC in later experiments.
  • TD3 (Fujimoto et al., 2018). A modified version of DDPG (Lillicrap, 2015) that addresses function approximation errors using three techniques: Clipped Double Q-Learning, Delayed Policy Updates, and Target Policy Smoothing.
  • TD3-BC (Fujimoto and Gu, 2021). An offline reinforcement learning variant of the TD3 algorithm, with behavior cloning (BC) regularization to constrain policy learning.
  • IQL (Kostrikov et al., 2021). An offline reinforcement learning algorithm that leverages expectile regression in Q-Learning.
Note that in our optimal live-stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique into TD3 and TD3-BC.
4.2.2. Variations of our model
We also compare the SL-MGAC algorithm with several variants in an ablation study to illustrate the effectiveness of multi-group state decomposition, supervised learning for critic learning, distribution discretization (DD), and other techniques.
  • SL-MGAC (w/o MG): An SL-MGAC variant without the Multi-Group State Decomposition module.
  • SL-MGAC (w/o MG & DD): An SL-MGAC variant without the Multi-Group State Decomposition module and the distribution discretization technique in reward learning.
  • SL-MGAC (w/o MG & DD & SL): An SL-MGAC variant without the Multi-Group State Decomposition module, the distribution discretization technique in reward learning, and the supervised learning procedure.
  • SL-MGAC (w/o LN): An SL-MGAC variant without the Layer Normalization technique.
  • SL-MGAC (w/o SG): An SL-MGAC variant without the Stop Gradient technique for hidden state representation in the actor network.
  • SL-MGAC (w/o Q-norm): An SL-MGAC variant without Q value normalization in the actor loss.
  • SL-MGAC-0: An SL-MGAC variant with the discount factor $\gamma=0$.
  • SL-MGAC-sep: An SL-MGAC variant with separate optimization of the actor network. Hence, $\hat{Q}_{norm}$ in Eq. LABEL:eq:15 is detached from the computation graph.
  • SL-MGAC-vanilla: An SL-MGAC variant with the vanilla Q label in the critic loss, as shown below, where $p(s_{t+1},\cdot)$ is the action probability of $s_{t+1}$ output by the actor network.
    (15) $Q_{label}(s_{t+1})$ $=r(s_{t},a_{t})+\gamma\sum_{a_{t+1}}p(s_{t+1},a_{t+1})\hat{Q}(s_{t+1},a_{t+1})$
4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with an embedding layer of size 5000. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods. The detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of Appendix B.
4.4. Offline Policy Evaluation
We follow the approach in (Xue et al., 2023) and adopt a commonly used offline policy evaluation method, namely Normalized Capped Importance Sampling (NCIS) (Swaminathan and Joachims, 2015), to evaluate performance. The evaluation metric is the cumulative reward across all trajectories of test users. Note that the reward for each sample is defined in Eq. 3.
Methods Cumulative Reward
L2R (Chen and Zhou, 2020) 413.49
DQN (Mnih et al., 2015) 413.18
BCQ (Fujimoto et al., 2019) 416.69
SAC (Christodoulou, [n. d.]) 435.41
TD3 (Fujimoto et al., 2018) 432.08
TD3-BC (Fujimoto and Gu, 2021) 430.23
IQL (Kostrikov et al., 2021) 432.03
SL-MGAC-sep 435.19
SL-MGAC-vanilla 450.21
SL-MGAC 458.49
Table 1. Overall offline performance of compared methods.
The offline policy evaluation results are shown in Table 1. Compared to the supervised learning method L2R, most RL methods achieve higher cumulative rewards, except for DQN. SL-MGAC significantly outperforms all other methods, demonstrating that SL-MGAC can achieve higher long-term rewards in the complex live-stream & short video mixed recommendation system with drastically changing data distributions. Furthermore, SL-MGAC outperforms SL-MGAC-vanilla, suggesting that incorporating the actor network in the Q label calculation of Eq. 15 may interfere with critic learning to some extent, slightly affecting the performance of SL-MGAC-vanilla. Additionally, SL-MGAC converges faster than SL-MGAC-sep during training, as shown in Fig. 8 in Appendix C, and achieves a higher cumulative reward in offline evaluation. This may be due to the information loss caused by detaching $\hat{Q}_{norm}$, which can lead to slower convergence for SL-MGAC-sep.
4.5. Ablation Study
We compare the offline performance of different SL-MGAC variants. As shown in Table 2, SL-MGAC outperforms all other variants on our offline dataset, demonstrating the effectiveness of the proposed multi-group state decomposition module, supervised learning procedure, and other techniques. Moreover, we find that SL-MGAC (w/o Q-norm) achieves the lowest reward, indicating that the Q-normalization technique in the actor loss enhances the model’s convergence and improves its performance.
Methods Cumulative Reward
SL-MGAC (w/o MG) 454.62
SL-MGAC (w/o MG & DD) 452.59
SL-MGAC (w/o MG & DD & SL) 449.13
SL-MGAC (w/o LN) 453.61
SL-MGAC (w/o SG) 457.26
SL-MGAC (w/o Q-norm) 392.12
SL-MGAC 458.49
Table 2. Offline performance of SL-MGAC variants.
Next, we compare the training processes of SL-MGAC and SL-MGAC (w/o MG), as shown in Fig. 5. We observe that the Q-value curve of SL-MGAC is more stable than that of SL-MGAC (w/o MG), and the Q-value variance of SL-MGAC is much smaller. This demonstrates the effectiveness and variance reduction effect of the MG-SD module.
Figure 5. The Q value curves between SL-MGAC and SL-MGAC (w/o MG) over 10 rounds of training. The lines correspond to the means of Q-value and the shaded areas correspond to the standard deviations.
4.6. Parameter Sensitivity
We analyze the impact of the number of user groups $K$ on the performance of SL-MGAC. As shown in Fig. 6, the proposed SL-MGAC algorithm with $K=6$ achieves the highest cumulative reward, demonstrating that increasing the number of user groups within a certain range improves model performance.
Figure 6. Performance of different numbers of user group $K$.
4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of live-stream daily active users (DAU) and watch time relative to the baseline. The baseline uses the SAC framework for online live-stream injection. The online A/B test results are shown in Table 3. SL-MGAC achieves the largest improvement in live-stream watch time while also increasing live-stream DAU. Although SL-MGAC-0 shows a larger improvement in live-stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live-streams, which may negatively impact the long-term user experience for most users. Therefore, this demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live-streams in response to user requests.
Methods Live-stream DAU Live-stream Watch Time
L2R +0.321% -0.558%
SL-MGAC-0 +3.928% -2.934%
SL-MGAC +2.616% +7.431%
Table 3. Online A/B Test Performance of compared methods.
Since the scale of live-stream supply and user interaction behaviors change drastically over time, more streamers tend to broadcast live during peak hours. As shown in Fig. 9 in Appendix C, SL-MGAC can adaptively adjust the live-stream injection strategy to increase live-stream exposure during peak hours, while reducing exposure during low-traffic periods.
4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 7, the live-stream injection ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live-stream injection ratio refers to the proportion of requests with action $a=1$.
Figure 7. The trend curve of online live-stream injection ratio during a whole day.
We calculate the amplitudes of the live-stream injection ratio within sliding time windows of 20 minutes. The amplitude density is shown in Fig. 10 in Appendix C. The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers. In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, which demonstrates the superior online model stability of SL-MGAC.
5. Related Work
5.1. RL in Recommendation Systems
Reinforcement learning (RL) aims to optimize long-term cumulative rewards over time, which has attracted significant attention in recommendation systems research in recent years (Afsar et al., 2022). Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend entire item lists, where the number of candidate items can be large. BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation. The work in (Miao et al., 2021) explores the use of off-policy RL for multi-session personalized search ranking. CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving user experience. Additionally, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
A similar approach, called self-supervised actor-critic (Xin et al., 2020), combines supervised learning with critic learning. Their supervised learning task focuses on predicting the next item recommendation probability, whereas our approach introduces supervised learning to restrict critic learning.
5.2. RL divergence and instability
Recently, there has been growing literature focusing on RL divergence and instability. (Van Hasselt et al., 2016) introduces Double DQN to mitigate the maximization bias problem in DQN. (Fujimoto et al., 2018) proposes the Clipped Double Q-Learning technique in TD3 to reduce the overestimation of Q-values. (Mao et al., 2018) introduces a bias-free, input-dependent baseline for the policy gradient algorithm (Sutton et al., 1999) to reduce variance and improve training stability. (Dohare et al., 2023) investigates policy collapse in a 2-state MDP and finds that L2 regularization and the non-stationary Adam optimizer (Kingma, 2014) are both effective in alleviating RL instability. (Yue et al., 2024) theoretically analyzes the causes of RL divergence and applies Layer Normalization to mitigate RL divergence and instability.
6. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live-stream injection policies in the challenging context of a live-stream & short video mixed recommendation system. Unlike existing RL methods for recommendation systems, our approach focuses on enhancing the stability and robustness of the RL model. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. In practice, we aim to minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully deployed in large-scale industrial recommendation systems.
7. Acknowledgment
We acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.
Appendix A Proof of Theorem 3.1
Proof.
Let state space $\mathcal{S}=\bigcup_{i=1}^{K}\mathcal{S}_{i},\forall i\neq j\;\mathcal{S}_{i}\cap\mathcal{S}_{j}=\varnothing$, state $s\in\mathcal{S}$. Let $r$ be the real reward of state $s$ and action $a$, $\hat{r}$ be the predicted value of $r$ over $\mathcal{S}$, $\hat{r}_{i}$ be the predicted reward of $r$ over $\mathcal{S}_{i}$.
Since a user has at most one request at timestamp $t$, a user will have at most one state $s$. We assume the state count $N$ of $\mathcal{S}$ is finite, the state count of $\mathcal{S}_{i}$ is $N_{i}$. Let $\mu$ be the mean value of $\hat{r}$ over $\mathcal{S}$, $\mu_{i}$ be the mean value of $\hat{r}_{i}$ over $\mathcal{S}_{i}$, then we have:
(16) $Var[\hat{r}]$ $=\frac{1}{N}\sum_{j=1}^{N}\left(\hat{r}_{j}-\mu\right)^{2}$
$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu\right)^{2}$
$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}+\mu_{i}-\mu\right)^{2}$
$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}\right)^{2}+\frac{1}{N}\sum_{i=1}^{K}N_{i}\left(\mu_{i}-\mu\right)^{2}$
where the first term is within-class variance $Var_{a}[\hat{r}]$ and the second term is between-class variance $Var_{b}[\hat{r}]$. As $Var_{b}[\hat{r}]>0$, we have
(17) $Var[\hat{r}]>\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}\right)^{2}=Var_{a}[\hat{r}]\$
As we partition the state space $\mathcal{S}$ into $K$ disjoint sub-spaces $\mathcal{S}_{i}$, then the variance of a neural network $f$ with multi-group state decomposition(MG SD) module is $Var_{a}[\hat{r}]$ while the variance of a neural network $f^{\prime}$ without MG-SD module is $Var[\hat{r}]$. ∎
Appendix B Tables
Hyper-parameter Value
MGN in Actor [128, 63, 31, 2]
MGN in RPN [128, 64, 32, 8]
MGN in QRN [128, 64, 32, 2]
Live-Time DDB [0s, 6s, 15s, 30s, 60s, 100s, 600s, 1200s]
Video-Time DDB [0s, 3s, 10s, 25s, 50s, 100s, 600s, 1200s]
Optimizer Adam
User Group Number $K$ 6
Online Exploration $\epsilon$ 0.2
Table 4. Hyper-parameters of SL-MGAC.
We denote MGN as the multi-group network, RPN as the Reward Prediction Network, QRN as Q Residual Networks, DDB as Distribution Discretization Bins. Both live and video watch time distribution have 7 time bins. We leave a separate time bin for the $a=1$ case and hence the output layer dimension of RPN is 8.
Appendix C Figures
Figure 8. Training Curves of SL-MGAC and SL-MGAC-sep.
Figure 9. Adaptive adjustments of live-stream exposure during peak hours and low-traffic periods.
Figure 10. The amplitude distribution of live-stream injection ratio.

### v2

@@
-Supervised Learning-enhanced Multi-Group Actor Critic for Live-stream Recommendation
+Supervised Learning-enhanced Multi-Group Actor Critic for Live Stream Allocation in Feed
 Abstract.
-Reinforcement Learning (RL) has been widely applied in recommendation systems to capture users’ long-term engagement, thereby improving dwelling time and enhancing user retention.
-In the context of a short video & live-stream mixed recommendation scenario, the live-stream recommendation system (RS) decides whether to inject at most one live-stream into the video feed for each user request.
-To maximize long-term user engagement, it is crucial to determine an optimal live-stream injection policy for accurate live-stream allocation.
-However, traditional RL algorithms often face divergence and instability problems, and these issues are even more pronounced in our scenario.
+Reinforcement Learning (RL) has been widely applied in recommendation systems to capture long-term user engagement, thus improving dwelling time and improving user retention.
+In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to inject at most one live stream into the video feed for each user request.
+To maximize long-term user engagement, it is crucial to determine an optimal live stream injection policy for accurate live stream allocation.
+However, traditional RL algorithms often face divergence and instability problems, and these issues may cause too many live stream allocation, which interrupts user’s short video interest and leads to a decrease in the user’s app usage duration.
 To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC).
@@
 Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability.
+We also propose a novel reward function to prevent overly greedy live stream allocation.
 Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing.
@@
 copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
-Figure 1. Structure of a short video & live-stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live-stream RS.
+Figure 1. Structure of a short video & live stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live stream RS.
 1. Introduction
-Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards across various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017).
+Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017).
 In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023).
 Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
-We consider a challenging sequential decision-making scenario in a short video & live-stream blended recommendation system, as shown in Fig.
+We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig.
 1.
-The system consists of three components: a live-stream recommendation system, a short video recommendation system, and a blending server.
-For each user request at timestamp $t$, the live-stream recommendation system decides whether to inject the recommended live-stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos.
-The blending server then mixes and rearranges the single live-stream with the $B$ videos to form the final recommendation list.
-Our objective is to find a personalized optimal live-stream injection policy that maximizes users’ long-term engagement with live-streams, which can be modeled as an infinite request-level Markov Decision Process (MDP) (Puterman, 1990).
-However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning (RL) algorithms (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario.
-A primary possible reason for this is the drastic fluctuations in both the live-stream supply scale and user interaction behaviors over time.
-As shown in Fig.
-2a, the curves of live-stream room count and live-stream viewer count vary significantly on a daily basis, making it difficult for the RL algorithm to learn optimal policies.
-Furthermore, as illustrated in Fig.
-2b, the distributions of user interaction labels (e.g., live-stream and short video watch time) are highly noisy and exhibit large variance.
-RL models often struggle to learn effective policies from data that exhibit such high variance across both time and user scales.
+The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server.
+For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos.
+The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list.
+Our objective is to find a personalized optimal live stream injection policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
+However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario.
+A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig.
+7 in Appendix B.
+RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales.
 In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
-Figure 2. Data distributions of online live-stream room (or viewer) count, live-stream watch time and short video watch time.
+Furthermore, like advertisement, live stream is not welcomed by every user.
+Allocating too many live streams for a user will significantly interrupt the user’s short video interest and lead to a decrease in the user’s app usage duration, which will eventually affect the user engagement and user retention.
+Therefore, we cannot merely focus on the user feedback for a single request, but rather should optimize the long-term user experience.
 To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC.
-Given the significant variation in users’ interest in live-stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks.
+Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate multi-group state decomposition (MG-SD) modules within both the actor and critic networks.
 Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation.
@@
   • We introduce SL-MGAC, a novel RL model designed to reduce prediction variance while enhancing the stability of both offline training and online inference.
-  • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q-value normalization.
+  • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q value normalization.
   • We integrate multi-task supervised learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
-  • We successfully deploy the SL-MGAC model in a challenging short video and live-stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
+  • We successfully deploy the SL-MGAC model in a challenging short video and live stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
 2. Problem Formulation
 As shown in Fig.
-1, in the short video & live-stream mixed recommendation system, the live-stream recommendation system is treated as an agent that interacts with diverse users and receives user feedback over time.
-The optimal live-stream injection control problem is modeled as an infinite request-level MDP to maximize the cumulative reward.
-Formally, we define the live-stream injection MDP as a tuple of five elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right)$:
-  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live-stream watch history, short video watch history), and item features (e.g., live-stream ID, author ID, author gender, online viewer count).
-We limit the length of users’ live-stream and short video history lists by timestamp order, keeping only the top 50 items.
-  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live-stream in response to a user’s request.
-We define action $a$ as a binary variable, where $a=1$ means injecting a live-stream, and $a=0$ means not injecting it.
+1, in the short video & live stream mixed recommendation system, the live stream recommendation system is treated as an agent that interacts with various users and receives user feedback over time.
+However, an agent considering instant user feedback of a single request may allocate live streams greedily in a short video feed and eventually affect user retention.
+Therefore, the optimal live stream injection control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
+Formally, we define live stream injection CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
+  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, online viewer count).
+We limit the length of user live stream and short video history lists by timestamp order, keeping only the top 50 items.
+  • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live stream in response to a user’s request.
+We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream and $a=0$ means not injecting it.
   • Transition Probability $\mathcal{P}$: The transition probability is denoted as $p(s_{t+1}|s_{t},a_{t})$, determined by the environment.
   • Reward Function $\mathcal{R}$: The reward function $\mathcal{R}$ is a mapping from state $s_{t}$ and action $a_{t}$ at timestamp $t$, which can be formulated as $r(s_{t},a_{t}):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$.
-  • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the computation of cumulative reward. Typically, we set $\gamma<1$ during model training.
+  • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the calculation of the cumulative reward. Typically, we set $\gamma<1$ during model training.
+  • Constraint $\mathcal{C}$: The platform-level constraint is that the live stream watch time should be longer than the average video watch time per request as much as possible.
+This prevents a negative impact on app usage duration, which could happen if the live stream watch time is shorter.
+By meeting this constraint, we ensure that live streams boost users’ long-term app engagement instead of reducing their time spent on it.
 Figure 2. Overall framework of the SL-MGAC algorithm. The SL (RL) MG-SD Module is short for the Multi-Group State Decomposition Module for supervised reward learning and critic learning.
-Overall, the optimization objective of aforementioned optimal live-stream injection control problem is shown as follows:
-(1) $\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
-where $\pi(a_{t}|s_{t})$ is the live-stream injection policy to be optimized.
-2.1. Reward Function Design
-In our mixed short video & live-stream recommendation scenario, we design a reward function based on the time gain of live-stream watch time relative to the average short video watch time in a request:
-(2) $r(s_{t},a_{t})=\begin{cases}y_{l}-\frac{1}{B}y_{v},\quad&if\;a_{t}=1\\ \beta&otherwise\\ \end{cases}$
-where $y_{l}$ is the live-stream watch time, $y_{v}$ is the total video watch time, $B$ is the number of videos in a request, and $\beta$ is the hyper-parameter in Eq.
+Overall, the optimization objective of aforementioned optimal live stream injection control problem is shown as follows:
+(1) $\max_{\pi}J_{R}(\pi)$
+$s.t.\quad J_{C}(\pi)\leq\epsilon$
+$J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
+$J_{C}(\pi)=\frac{1}{B}y_{v}-y_{l}$
+where $\pi(a_{t}|s_{t})$ is the live stream injection policy to be optimized, $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
+A common method to solve the above CMDP is to transform the problem into a min-max optimization by introducing the Lagrange multiplier $\lambda$:
+(2) $(\pi^{*},\lambda^{*})=\arg\min_{\lambda\geq 0}\max_{\pi}J_{R}(\pi)-\lambda(J_{C}(\pi)-\epsilon)$
+Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function:
+(3) $\hat{r}(\lambda,s_{t},a_{t})$ $=r(s_{t},a_{t})-\lambda c(s_{t},a_{t})$
+$=y_{l}-\lambda(\frac{1}{B}y_{v}-y_{l})$
+$=(1+\lambda)y_{l}-\frac{\lambda}{B}y_{v}$
+In practice, we solve the unconstrained MDP problem and the reward function is simplified as follows:
+(4) $r(s_{t},a_{t})=y_{l}-\frac{\lambda}{B}y_{v}$
+where $\lambda$ is a hyper-parameter to carefully control the final live stream injection ratio.
+From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint.
+3. Proposed Framework
+To address the intrinsic divergence and instability issues of reinforcement learning and successfully deploy the RL agent in our mixed short video & live stream recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC), as shown in Fig.
 2.
-Note that the recommended list in a request always contains $B$ videos and at most one live-stream.
-Furthermore, the reward defined above is similar to that of the RCPO (Tessler et al., 2018) algorithm.
-From the perspective of constrained reinforcement learning, the term $\frac{1}{B}y_{v}$ can be viewed as an implicit cost constraint.
-In practice, the hyper-parameter $\beta$ is introduced to carefully control the final live-stream injection ratio.
-Since the distribution of live-stream (short video) watch time changes drastically over time, as shown in Fig.
-2b, it is challenging for an RL agent to optimize the cumulative reward in Eq.
-2.
-To address this, we propose a reward normalization technique as follows:
-(3) $r(s_{t},a_{t})=\begin{cases}sigmoid\left(y_{l}-\frac{1}{B}y_{v}\right),\quad&if\;a_{t}=1\\ \beta\in[0,1]&otherwise\\ \end{cases}$
-which helps stabilize the network gradient, thereby reducing the risk of model divergence.
-3. Proposed Framework
-To address the intrinsic divergence and instability issues of reinforcement learning and successfully deploy the RL agent in our mixed short video & live-stream recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC), as shown in Fig.
-3.
-SL-MGAC incorporates a supervised learning-enhanced actor-critic model with a shared user & live-stream feature extraction module, along with independent Multi-Group State Decomposition (MG-SD) modules.
+SL-MGAC incorporates a supervised learning-enhanced actor-critic model with a shared user & live stream feature extraction module, along with independent Multi-Group State Decomposition (MG-SD) modules.
 For clarity, only one critic network is shown in Fig.
-3, while the remaining three critic networks, which use the same architecture, are omitted.
-3.1. User & Live-stream Feature Extraction Module
-The user & live-stream feature extraction module aims to generate non-linear embedding representations for the state $s_{t}$.
-First, we use a unified embedding layer to map the user’s static features, live-stream features, historical live-stream list features and short video list features into low-dimensional dense representations.
+2, while the remaining three critic networks, which use the same architecture, are omitted.
+3.1. User & live stream Feature Extraction Module
+The user & live stream feature extraction module aims to generate non-linear embedding representations for the state $s_{t}$.
+First, we use a unified embedding layer to map the user’s static features, live stream features, historical live stream list features, and short video-list features into low-dimensional dense representations.
 We denote $\mathbf{v}_{U}$, $\mathbf{v}_{L}$, $\left\{\mathbf{e}_{1},\cdots,\mathbf{e}_{M_{l}}\right\}$, $\left\{\mathbf{e}^{\prime}_{1},\cdots,\mathbf{e}^{\prime}_{M_{v}}\right\}$ as the corresponding embedding vectors or sets of embedding vectors, respectively.
 We then define $\mathbf{h}_{a}=[\mathbf{v}_{U},\mathbf{v}_{L}]$ as the concatenation of $\mathbf{v}_{U}$ and $\mathbf{v}_{L}$.
-To aggregate historical live-stream (short video) representations, we introduce the target attention mechanism (Vaswani, 2017; Zhou et al., 2018), which is defined as follows:
+To aggregate historical live stream (short video) representations, we introduce the target attention mechanism (Vaswani, 2017; Zhou et al., 2018), which is defined as follows:
 (5) $\mathbf{h}_{live}$ $=\sum_{i=0}^{M_{l}}f_{l}(\mathbf{h}_{a},\mathbf{e}_{i})\mathbf{e}_{i}$
@@
 where $f_{l}$, $f_{v}$ are different target attention functions, such as a feed-forward network whose output is a normalized score.
-After aggregating the historical live-stream and short video features via two separate attention networks, we concatenate all the above embedding vectors to form $\mathbf{h}_{s}=[\mathbf{v}_{U},\mathbf{v}_{L},\mathbf{h}_{live},\mathbf{h}_{video}]$.
+After aggregating the historical live stream and short video features via two separate attention networks, we concatenate all of the above embedding vectors to form $\mathbf{h}_{s}=[\mathbf{v}_{U},\mathbf{v}_{L},\mathbf{h}_{live},\mathbf{h}_{video}]$.
 We then use a shared multi-layer perceptron (MLP) for both the actor and critic networks to obtain the latent representation of state $s_{t}$, i.e.
@@
 3.2. Multi-Group State Decomposition Module
-Since live-stream content is not equally appealing to all users on the short video app Kwai, it is crucial to inject live-streams selectively into the short video feed.
-Otherwise, excessive live-stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration.
-In practice, user interaction data for live-streams is sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live-stream injection policy for each individual user.
-Therefore, prior information about diverse user groups is essential for enhancing the decision-making accuracy of RL models.
+Since live stream content is not equally appealing to all users of the short video app Kwai, it is crucial to inject live streams selectively into the short video feed.
+Otherwise, excessive live stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration.
+In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream injection policy for each individual user.
+Directly learning a policy from sparse and noisy user feedback is often infeasible, since it will inadvertently inject excessive live streams in the short video feed and highly affected the user engagement.
+We find that prior information on diverse user groups is essential to improve the accuracy of the decision making of RL models in our scenario, which provides additional user preference information to the policy.
 A natural approach to distinguish between different user behaviors is to partition users into distinct groups based on their historical activity.
-Specifically, we categorize users into $K$ disjoint groups according to their live-stream activity level, which is determined by the cumulative live-stream watch time over the past 3 weeks.
-Users with higher historical watch time are assigned to groups with higher activity levels.
-Formally, we present the following theorem, which demonstrates that the proposed multi-group state decomposition can be viewed as a type of variance reduction technique for real-world RL applications.
-Theorem 3.1.
-Let state space $\mathcal{S}=\bigcup_{i=1}^{K}\mathcal{S}_{i},\forall i\neq j\;\mathcal{S}_{i}\cap\mathcal{S}_{j}=\varnothing$, state $s\in\mathcal{S}$; $r$ be the real reward of state $s$ and action $a$, $\hat{r}$ be the predicted value of $r$ over $\mathcal{S}$, $\hat{r}_{i}$ be the predicted reward of $r$ over $\mathcal{S}_{i}$; then a neural network $f$ with MG-SD module has a lower prediction variance than the neural network $f^{\prime}$ without MG-SD module.
-Please refer to Appendix.
-A for detailed proof of Theorem 3.1.
-We demonstrate that the MG-SD module eliminates between-group variance, thereby reducing the total variance, which implies that the multi-group state decomposition module contributes to a more stable RL model.
-Empirically, we also show the effectiveness of this module in enhancing RL model stability in subsequent experiments.
+Specifically, we categorize users into $K$ disjoint groups according to their activity level of the live stream, which is determined by the cumulative watch time of the live stream during the past 3 weeks.
+Users with a higher historical watch time are assigned to groups with higher activity levels.
+Empirically, we show the effectiveness of this module in enhancing the stability of the RL model in subsequent experiments.
 We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
@@
 Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse.
-Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate the divergence and instability problems.
-From the proof in the Appendix D of (Yue et al., 2024), we know that for any input $\mathbf{x}$ and any direction $\mathbf{v}$, if a network applies Layer Normalization to the input, then we have $k_{\rm NTK}(\mathbf{x},\mathbf{x}+\lambda\mathbf{v})\leq C$, where $\lambda>0$, $C$ is a constant and $k_{\rm NTK}$ is the Neural Tangent Kernel (NTK) (Jacot et al., 2018).
+Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate divergence and instability problems.
+From the proof in Appendix D of (Yue et al., 2024), we know that for any input $\mathbf{x}$ and any direction $\mathbf{v}$, if a network applies Layer Normalization to the input, then we have $k_{\rm NTK}(\mathbf{x},\mathbf{x}+\lambda\mathbf{v})\leq C$, where $\lambda>0$, $C$ is a constant, and $k_{\rm NTK}$ is the Neural Tangent Kernel (NTK) (Jacot et al., 2018).
 This theoretically indicates that a network with Layer Normalization is less sensitive to input variations and can maintain stable gradients despite perturbations during model training.
 Hence, we apply Layer Normalization to the inputs of both the actor and critic networks, as shown in Fig.
-3.
+2.
 3.3.2. Critic Network
@@
 where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use Huber loss (Hastie et al., 2009) to optimize the above objective.
-In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of the critic networks.
+In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of critic networks.
+Specifically, due to the instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term.
+Then the $r(s_{t},a_{t})$ term could not provide any information to guide the critic learning.
 Therefore, we seamlessly introduce supervised learning (e.g.
 multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure.
-Specifically, we split the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
+Specifically, we divide the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
 (7) $Q_{\phi_{i}}(s_{t},a_{t})$ $=R_{\theta_{i}}(s_{t},a_{t})+\gamma T_{\xi_{i}}(s_{t},a_{t}),\quad i=1,2$
@@
 where $R_{\theta_{i}}$ and $T_{\xi_{i}}$ are RPN and QRN, $R^{\prime}_{\theta_{i}}$ and $T^{\prime}_{\xi_{i}}$ are target RPN and QRN, respectively. We use separate MLPs to model the RPN and QRN.
-Since the distribution of live-stream (short video) watch time changes drastically over time, the time gain reward $r$ in Eq.
+Since the distribution of live stream (short video) watch time changes drastically over time, the time gain reward $r$ in Eq.
 4 becomes difficult to learn.
 In this work, we propose a distribution discretization method to improve reward learning.
-Formally, we divide the live-stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live-stream (short video) watch time.
-Let $y$ denote the actual live-stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$.
+Formally, we divide the live stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live stream (short video) watch time.
+Let $y$ denote the actual live stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$.
 The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$.
@@
 which resembles the structure of linear reward shifting (Sun et al., 2022).
-Let $o(t)_{l}\in\mathbb{R}^{N_{l}+1}$, $o(t)_{v}\in\mathbb{R}^{N_{v}+1}$ be one-hot vectors representing the time bins in which the real live-stream (short video) watch time of sample $(s_{t},a_{t},r_{t})$ falls.
+Let $o(t)_{l}\in\mathbb{R}^{N_{l}+1}$, $o(t)_{v}\in\mathbb{R}^{N_{v}+1}$ be one-hot vectors representing the time bins in which the real live stream (short video) watch time of sample $(s_{t},a_{t},r_{t})$ falls.
 Note that a separate bin is set for the case $a_{t}=0$.
@@
 $G_{\Theta_{i}}(s_{t},a_{t})$ $={o(t)_{v}}^{T}\left(y_{v}^{st}+g_{\Theta_{i}}(s_{t},a_{t})\cdot\left(y_{v}^{end}-y_{v}^{st}\right)\right)$
-where $y_{l}^{st},y_{l}^{end}\in\mathbb{R}^{N_{l}+1}$ predefined left (right) boundary vectors for the live-stream watch time bin, $y_{v}^{st},y_{v}^{end}\in\mathbb{R}^{N_{v}+1}$ are predefined left (right) boundary vectors for the short video watch time bin.
-Note that introducing the posterior one-hot vectors $o(t)_{l}$ and $o(t)_{v}$ in Eq.
+where $y_{l}^{st},y_{l}^{end}\in\mathbb{R}^{N_{l}+1}$ predefined left (right) boundary vectors for the live stream watch time bin, $y_{v}^{st},y_{v}^{end}\in\mathbb{R}^{N_{v}+1}$ are predefined left (right) boundary vectors for the short video watch time bin.
+Note that we introduce the posterior one-hot vectors $o(t)_{l}$ and $o(t)_{v}$ in Eq.
 9 will not affect the calculation of $Q^{\prime}_{\phi_{i}}(s_{t+1},\cdot)$, because we have already reserved $r_{t+1}$ in the dataset $\mathcal{D}$.
@@
 As shown in Eq.
-8, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin.
-We then reconstruct the predicted live-stream and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively.
+9, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin.
+We then reconstruct the predicted live streaming and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively.
 Finally, we obtain the predicted normalized reward, $R_{\theta_{i}}(s_{t},a_{t})$.
-To optimize the predicted watch time proportions, we use Huber Loss as follows:
+To optimize the predicted watch-time proportions, we use Huber loss as follows:
 (10) $\mathcal{L}_{SL}=\sum_{i=1}^{2}Huber\_Loss(\delta_{l},F^{\prime}_{\Gamma_{i}})+Huber\_Loss(\delta_{v},G^{\prime}_{\Theta_{i}})$
-where $\delta_{l}$, $\delta_{v}$ are the actual proportion labels for live-stream and short video watch times, respectively.
-The reason for using Huber loss on the time proportion $\delta$ instead of the actual watch time $y$ is that the time proportion lies within the range $[0,1]$, which results in smaller gradients in the neural network.
-Additionally, the variance of $\delta$ is much smaller than that of the original watch time $y$.
+where $\delta_{l}$, $\delta_{v}$ are the actual proportion labels for live stream and short video watch times, respectively.
+The reason for using the Huber loss on the time ratio $\delta$ instead of the actual watch time $y$ is that the time ratio is within the range $[0,1]$, resulting in smaller gradients in the neural network.
+Furthermore, the variance of $\delta$ is much smaller than that of the original watch time $y$.
 Consequently, the outputs of the learned reward networks ($f_{\Gamma_{i}}$ and $g_{\Theta_{i}}$) will exhibit smaller variances.
@@
 where $\hat{Q}(s_{t},a_{t})=\min_{i=1,2}Q_{\phi_{i}}(s_{t},a_{t})$, and $p(s_{t},a_{t})$ is the action probability output of the actor network.
-Moreover, we observe that large values of $\hat{Q}(s_{t},a_{t})$ may cause instability in actor training, which can ultimately lead to policy deterioration, where $\forall t,\;\pi(a_{t}=1|s_{t})>\pi(a_{t}=0|s_{t})$ or $\pi(a_{t}=1|s_{t})<\pi(a_{t}=0|s_{t})$.
+Moreover, we observe that high values of $\hat{Q}(s_{t},a_{t})$ can cause instability in actor training, which can ultimately lead to policy deterioration, where $\forall t,\;\pi(a_{t}=1|s_{t})>\pi(a_{t}=0|s_{t})$ or $\pi(a_{t}=1|s_{t})<\pi(a_{t}=0|s_{t})$.
 To address this, we apply softmax normalization to $\hat{Q}(s_{t},a_{t})$ to obtain $\hat{Q}_{norm}=softmax(\hat{Q}(s_{t},a_{t}))$, and propose a modified actor loss function:
@@
 The teacher critic networks guide the more lightweight student actor network to adjust and converge to an optimal policy.
-In practice, we only need to deploy the actor network in the online live-stream recommendation system, which significantly reduces computational complexity and improves real-time respond speed.
+In practice, we only need to deploy the actor network in the online live stream recommendation system, which significantly reduces computational complexity and improves real-time response speed.
 Overall, the final loss function of our proposed SL-MGAC algorithm is as follows:
@@
 where $\psi$ is a random number, and $\epsilon$ is maximal exploration probability.
-We deploy the SL-MGAC algorithm in our recommendation system, and the overall system architecture is shown in Fig.
-4.
+We implemented the SL-MGAC algorithm in our recommendation system, and the overall architecture of the system is shown in Fig.
+3.
 The online RL agent collects real-time user interaction logs, while the offline model trainer optimizes the SL-MGAC model in an off-policy manner using streaming training data.
@@
 4. Experiments
-We conduct both offline evaluation on a real-world dataset collected from our recommendation system and online A/B test experiments with SL-MGAC and the baselines.
+We perform both offline evaluation on a real-world dataset collected from our recommendation system and online A/B test experiments with SL-MGAC and the baselines.
 4.1. Dataset
-Due to the lack of publicly available datasets for recommendation decisions involving live-streaming injections in short video feeds, we collect a dataset from our recommendation system with 10,000 users through random sampling over a full day, ensuring that the proportions of different user groups are similar to those in real online data.
-In total, we have over 180,000 samples, which are split into training and test sets with a 4:1 ratio.
+Due to the lack of publicly available datasets for recommendation decisions involving live streaming injections in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data.
+In total, we have over 1,800,000 samples, which are divided into training and test sets with a 4:1 ratio.
+In addition, the online environment contains more than 100 million app users for AB-Tests.
 4.2. Compared Methods
@@
   • Learning to Rank (L2R) (Chen and Zhou, 2020). A supervised learning framework that predicts the reward for each action. The action with the maximum reward is selected as the agent’s action.
-  • DQN (Mnih et al., 2015). A deep neural network algorithm for Q-learning that introduces the technique of target network updates.
+  • DQN (Mnih et al., 2015). A deep neural network algorithm for Q-learning that introduces the technique of updating the target network.
   • BCQ (Fujimoto et al., 2019). A widely used offline reinforcement learning algorithm that adds action restrictions and encourages the agent to behave similarly to the behavior policy.
@@
   • TD3 (Fujimoto et al., 2018).
-A modified version of DDPG (Lillicrap, 2015) that addresses function approximation errors using three techniques: Clipped Double Q-Learning, Delayed Policy Updates, and Target Policy Smoothing.
+A modified version of DDPG (Lillicrap, 2015) addresses function approximation errors using three techniques: Clipped Double Q-Learning, Delayed Policy Updates, and Target Policy Smoothing.
   • TD3-BC (Fujimoto and Gu, 2021). An offline reinforcement learning variant of the TD3 algorithm, with behavior cloning (BC) regularization to constrain policy learning.
   • IQL (Kostrikov et al., 2021). An offline reinforcement learning algorithm that leverages expectile regression in Q-Learning.
-Note that in our optimal live-stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique into TD3 and TD3-BC.
+Note that in our optimal live stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
 4.2.2. Variations of our model
@@
   • SL-MGAC (w/o SG): An SL-MGAC variant without the Stop Gradient technique for hidden state representation in the actor network.
-  • SL-MGAC (w/o Q-norm): An SL-MGAC variant without Q value normalization in the actor loss.
+  • SL-MGAC (w/o Q-norm): An SL-MGAC variant without Q-value normalization in actor loss.
   • SL-MGAC-0: An SL-MGAC variant with the discount factor $\gamma=0$.
   • SL-MGAC-sep: An SL-MGAC variant with separate optimization of the actor network. Hence, $\hat{Q}_{norm}$ in Eq. LABEL:eq:15 is detached from the computation graph.
-  • SL-MGAC-vanilla: An SL-MGAC variant with the vanilla Q label in the critic loss, as shown below, where $p(s_{t+1},\cdot)$ is the action probability of $s_{t+1}$ output by the actor network.
+  • SL-MGAC-vanilla: An SL-MGAC variant with the vanilla Q label in the loss of the critic, as shown below, where $p(s_{t+1},\cdot)$ is the probability of action of the output of $s_{t+1}$ by the actor network.
     (16) $Q_{label}(s_{t+1})$ $=r(s_{t},a_{t})+\gamma\sum_{a_{t+1}}p(s_{t+1},a_{t+1})\hat{Q}(s_{t+1},a_{t+1})$
 4.3. Implementation Details
-To ensure fairness among all compared methods, we use a consistent feature extraction module with an embedding layer of size 5000.
+To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer.
 In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128].
@@
 The discount factor $\gamma$ is set to 0.9 for all methods.
-The detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of Appendix B.
+Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
 4.4. Offline Policy Evaluation
@@
 The evaluation metric is the cumulative reward across all trajectories of test users.
-Note that the reward for each sample is defined in Eq.
-3.
 Methods Cumulative Reward
@@
 Table 1. Overall offline performance of compared methods.
-The offline policy evaluation results are shown in Table 1.
+The results of the offline policy evaluation are shown in Table 1.
 Compared to the supervised learning method L2R, most RL methods achieve higher cumulative rewards, except for DQN.
-SL-MGAC significantly outperforms all other methods, demonstrating that SL-MGAC can achieve higher long-term rewards in the complex live-stream & short video mixed recommendation system with drastically changing data distributions.
+SL-MGAC significantly outperforms all other methods, demonstrating that SL-MGAC can achieve higher long-term rewards in the complex live stream & short video mixed recommendation system with drastically changing data distributions.
 Furthermore, SL-MGAC outperforms SL-MGAC-vanilla, suggesting that incorporating the actor network in the Q label calculation of Eq.
-15 may interfere with critic learning to some extent, slightly affecting the performance of SL-MGAC-vanilla.
-Additionally, SL-MGAC converges faster than SL-MGAC-sep during training, as shown in Fig.
-8 in Appendix C, and achieves a higher cumulative reward in offline evaluation.
-This may be due to the information loss caused by detaching $\hat{Q}_{norm}$, which can lead to slower convergence for SL-MGAC-sep.
+16 may interfere with the learning of the critic to some extent, slightly affecting the performance of the SL-MGAC-vanilla.
 4.5. Ablation Study
@@
 Next, we compare the training processes of SL-MGAC and SL-MGAC (w/o MG), as shown in Fig.
-5.
+4.
 We observe that the Q-value curve of SL-MGAC is more stable than that of SL-MGAC (w/o MG), and the Q-value variance of SL-MGAC is much smaller.
@@
 As shown in Fig.
-6, the proposed SL-MGAC algorithm with $K=6$ achieves the highest cumulative reward, demonstrating that increasing the number of user groups within a certain range improves model performance.
+5, the proposed SL-MGAC algorithm with $K=6$ achieves the highest cumulative reward, demonstrating that increasing the number of user groups within a certain range improves model performance.
 Figure 5. Performance of different numbers of user group $K$.
+Methods live stream DAU live stream Watch Time Video Watch Time App Usage Duration User Retention
+L2R +0.321% -0.558% -0.238% -0.178% -0.154%
+Dummy +51.3% +34.6% -2.216% -1.062% -0.439%
+SL-MGAC-0 +3.928% -2.934% -0.153% -0.117% -0.102%
+SL-MGAC +2.616% +7.431% +0.197% +0.121% +0.086%
+Table 3. Online A/B Test Performance of compared methods.
 4.7. Online A/B Test Experiment
-We compare the improvements of different methods in terms of live-stream daily active users (DAU) and watch time relative to the baseline.
-The baseline uses the SAC framework for online live-stream injection.
-The online A/B test results are shown in Table 3.
-SL-MGAC achieves the largest improvement in live-stream watch time while also increasing live-stream DAU.
-Although SL-MGAC-0 shows a larger improvement in live-stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live-streams, which may negatively impact the long-term user experience for most users.
-Therefore, this demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live-streams in response to user requests.
-Methods Live-stream DAU Live-stream Watch Time
-L2R +0.321% -0.558%
-SL-MGAC-0 +3.928% -2.934%
-SL-MGAC +2.616% +7.431%
-Table 3. Online A/B Test Performance of compared methods.
-Since the scale of live-stream supply and user interaction behaviors change drastically over time, more streamers tend to broadcast live during peak hours.
-As shown in Fig.
-9 in Appendix C, SL-MGAC can adaptively adjust the live-stream injection strategy to increase live-stream exposure during peak hours, while reducing exposure during low-traffic periods.
+We compare the improvements of different methods in terms of live stream daily active users (DAU), live stream watch time, video watch time, app usage duration and user retention relative to the baseline.
+The baseline uses the SAC framework for live stream injection.
+The results of the online A / B test are shown in Table 3.
+Note that the Dummy method injects a live stream for each request.
+We can see that SL-MGAC achieves the largest improvement in live stream watch time while also increasing the app usage duration and user retention than other methods.
+It demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live streams in response to user requests.
+Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may negatively impact the long-term user experience for most users.
+Moreover, the dummy method can greatly improve live streaming DAU and watch time, but significantly affect app usage time and user retention.
 4.8. Online Model Stability
@@
 As shown in Fig.
-7, the live-stream injection ratio fluctuates drastically throughout the day for both models.
+6, the live stream injection ratio fluctuates drastically throughout the day for both models.
 However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude.
-Note that the live-stream injection ratio refers to the proportion of requests with action $a=1$.
-Figure 7. The trend curve of online live-stream injection ratio during a whole day.
-We calculate the amplitudes of the live-stream injection ratio within sliding time windows of 20 minutes.
+Note that the live stream injection ratio refers to the proportion of requests with action $a=1$.
+Figure 6. The trend curve of online live stream injection ratio during a whole day.
+We calculate the amplitudes of the live stream injection ratio within sliding time windows of 20 minutes.
 The amplitude density is shown in Fig.
-10 in Appendix C.
+8 in Appendix B.
 The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers.
-In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, which demonstrates the superior online model stability of SL-MGAC.
+In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, demonstrating the superior online model stability of SL-MGAC.
+5. Discussion
+The intrinsic instability and divergence of RL limits its application in real-world industrial recommendation systems.
+The proposed SL-MGAC method improves the robustness and stability of the RL model and achieves significant offline and online performance in the live stream and short video mixed recommendation scenario.
+However, SL-MGAC only learns from data from adjacent states $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1},r_{t+1})$, which limits its ability to learn information in much longer sequences, such as the daily-level user preference, rather than session-level user preference, since the capture of daily user preference information is very helpful for improving user engagement and user retention.
+Learning live stream policy from a long sequence of user behaviors would lead to our future work.
+Furthermore, we emphasize that more sophisticated methods can also be employed to divide users into multiple groups and provide prior information about user preference, such as traditional clustering methods, e.g.
+k-means, neural clustering of VQ-VAE (Van Den Oord et al., 2017), online clustering of SwAV (Caron et al., 2020) by the Sinkhorn-Knopp algorithm (Cuturi, 2013), etc.
+We find that the simple MG-SD module achieves significant performance and fast real-time response spped in practice; we will lead these complicated group partitioning methods in our further research.
 6. Related Work
 6.1. RL in Recommendation Systems
-Reinforcement learning (RL) aims to optimize long-term cumulative rewards over time, which has attracted significant attention in recommendation systems research in recent years (Afsar et al., 2022).
-Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend entire item lists, where the number of candidate items can be large.
+Reinforcement learning (RL) aims to optimize cumulative long-term rewards over time, which has attracted significant attention in the research of recommendation systems in recent years (Afsar et al., 2022).
+Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend complete item lists, where the number of candidate items can be large.
 BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation.
-The work in (Miao et al., 2021) explores the use of off-policy RL for multi-session personalized search ranking.
-CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving user experience.
-Additionally, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
+The work in (Miao et al., 2021) explores the use of off-policy RL for personalized search ranking in multiple sessions.
+CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving the user experience.
+Furthermore, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
 A similar approach, called self-supervised actor-critic (Xin et al., 2020), combines supervised learning with critic learning.
@@
 6.2. RL divergence and instability
-Recently, there has been growing literature focusing on RL divergence and instability.
+Recently, there has been a growing literature that focuses on RL divergence and instability.
 (Van Hasselt et al., 2016) introduces Double DQN to mitigate the maximization bias problem in DQN.
-(Fujimoto et al., 2018) proposes the Clipped Double Q-Learning technique in TD3 to reduce the overestimation of Q-values.
-(Mao et al., 2018) introduces a bias-free, input-dependent baseline for the policy gradient algorithm (Sutton et al., 1999) to reduce variance and improve training stability.
+(Fujimoto et al., 2018) proposes the Clipped Double Q-Learning technique in TD3 to reduce the overestimation of Q values.
+(Mao et al., 2018) introduces a bias-free, input-dependent baseline for the policy gradient algorithm (Sutton et al., 1999) to reduce variance and improve the stability of training.
 (Dohare et al., 2023) investigates policy collapse in a 2-state MDP and finds that L2 regularization and the non-stationary Adam optimizer (Kingma, 2014) are both effective in alleviating RL instability.
@@
 7. Conclusion
-In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live-stream injection policies in the challenging context of a live-stream & short video mixed recommendation system.
+In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream injection policies in the challenging context of a live stream & short video mixed recommendation system.
 Unlike existing RL methods for recommendation systems, our approach focuses on enhancing the stability and robustness of the RL model.
@@
 In practice, we aim to minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully deployed in large-scale industrial recommendation systems.
-7. Acknowledgment
-We acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.
-Appendix A Proof of Theorem 3.1
-Proof.
-Let state space $\mathcal{S}=\bigcup_{i=1}^{K}\mathcal{S}_{i},\forall i\neq j\;\mathcal{S}_{i}\cap\mathcal{S}_{j}=\varnothing$, state $s\in\mathcal{S}$.
-Let $r$ be the real reward of state $s$ and action $a$, $\hat{r}$ be the predicted value of $r$ over $\mathcal{S}$, $\hat{r}_{i}$ be the predicted reward of $r$ over $\mathcal{S}_{i}$.
-Since a user has at most one request at timestamp $t$, a user will have at most one state $s$.
-We assume the state count $N$ of $\mathcal{S}$ is finite, the state count of $\mathcal{S}_{i}$ is $N_{i}$.
-Let $\mu$ be the mean value of $\hat{r}$ over $\mathcal{S}$, $\mu_{i}$ be the mean value of $\hat{r}_{i}$ over $\mathcal{S}_{i}$, then we have:
-(16) $Var[\hat{r}]$ $=\frac{1}{N}\sum_{j=1}^{N}\left(\hat{r}_{j}-\mu\right)^{2}$
-$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu\right)^{2}$
-$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}+\mu_{i}-\mu\right)^{2}$
-$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}\right)^{2}+\frac{1}{N}\sum_{i=1}^{K}N_{i}\left(\mu_{i}-\mu\right)^{2}$
-where the first term is within-class variance $Var_{a}[\hat{r}]$ and the second term is between-class variance $Var_{b}[\hat{r}]$. As $Var_{b}[\hat{r}]>0$, we have
-(17) $Var[\hat{r}]>\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}\right)^{2}=Var_{a}[\hat{r}]\$
-As we partition the state space $\mathcal{S}$ into $K$ disjoint sub-spaces $\mathcal{S}_{i}$, then the variance of a neural network $f$ with multi-group state decomposition(MG SD) module is $Var_{a}[\hat{r}]$ while the variance of a neural network $f^{\prime}$ without MG-SD module is $Var[\hat{r}]$.
-∎
-Appendix B Tables
+Appendix A Tables
 Hyper-parameter Value
@@
 Table 4. Hyper-parameters of SL-MGAC.
-We denote MGN as the multi-group network, RPN as the Reward Prediction Network, QRN as Q Residual Networks, DDB as Distribution Discretization Bins.
+We denote MGN as the multi-group network, RPN as the Reward Prediction Network, QRN as Q Residual Networks, and DDB as Distribution Discretization Bins.
 Both live and video watch time distribution have 7 time bins.
-We leave a separate time bin for the $a=1$ case and hence the output layer dimension of RPN is 8.
-Appendix C Figures
-Figure 8. Training Curves of SL-MGAC and SL-MGAC-sep.
-Figure 9. Adaptive adjustments of live-stream exposure during peak hours and low-traffic periods.
-Figure 10. The amplitude distribution of live-stream injection ratio.
+We leave a separate time bin for the $a=1$ case, and hence the output layer dimension of RPN is 8.
+Appendix B Figures
+Figure 7. Data distributions of online live stream room (or viewer) count, live stream watch time and short video watch time.
+Figure 8. The amplitude distribution of live stream injection ratio.

### v4

@@
 Abstract.
-Reinforcement Learning (RL) has been widely applied in recommendation systems to capture long-term user engagement, thus improving dwelling time and improving user retention.
-In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to inject at most one live stream into the video feed for each user request.
-To maximize long-term user engagement, it is crucial to determine an optimal live stream injection policy for accurate live stream allocation.
-However, traditional RL algorithms often face divergence and instability problems, and these issues may cause too many live stream allocation, which interrupts user’s short video interest and leads to a decrease in the user’s app usage duration.
+In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream into the video feed for each user request.
+To maximize long-term user engagement, it is crucial to determine an optimal live stream policy for accurate live stream allocation.
+The inappropriate live stream allocation policy can significantly affect the duration of the usage app and user retention, which ignores the long-term negative impact of live stream allocation.
+Recently, reinforcement learning (RL) has been widely applied in recommendation systems to capture long-term user engagement.
+However, traditional RL algorithms often face divergence and instability problems, which restricts the application and deployment in the large-scale industrial recommendation systems, especially in the aforementioned challenging scenario.
 To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC).
@@
 Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing.
-Experimental results demonstrate that the proposed method not only outperforms baseline methods but also exhibits enhanced stability in online recommendation scenarios.
+Experimental results demonstrate that the proposed method not only outperforms baseline methods under the platform-level constraints but also exhibits enhanced stability in online recommendation scenarios.
 Reinforcement Learning, Recommendation System, Variance Reduction
@@
 1. Introduction
-Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017).
-In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023).
-Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
 We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig.
@@
 The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list.
-Our objective is to find a personalized optimal live stream injection policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
-However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario.
+Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
+Currently, deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017).
+In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023).
+Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
+However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system.
 A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig.
@@
 Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
-The main contributions of this paper are as follows:
-  • We introduce SL-MGAC, a novel RL model designed to reduce prediction variance while enhancing the stability of both offline training and online inference.
+The main contributions of this paper are as follows.
+  • We introduce SL-MGAC, a novel RL model that outperforms existing methods in maximizing the user engagement of live stream while satisfying the platform-level constraints on app usage duration and user retention.
   • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q value normalization.
@@
 However, an agent considering instant user feedback of a single request may allocate live streams greedily in a short video feed and eventually affect user retention.
-Therefore, the optimal live stream injection control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
-Formally, we define live stream injection CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
+Therefore, the optimal live stream allocation control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
+Formally, we define live stream allocation CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
   • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, online viewer count).
@@
   • Discount Factor $\gamma$: The discount factor $\gamma\in[0,1]$ is used in the calculation of the cumulative reward. Typically, we set $\gamma<1$ during model training.
-  • Constraint $\mathcal{C}$: The platform-level constraint is that the live stream watch time should be longer than the average video watch time per request as much as possible.
+  • Constraint $\mathcal{C}$: The platform-level constraint is an instantaneous constraint, that is, the live stream watch time should be longer than the average video watch time per request as much as possible.
 This prevents a negative impact on app usage duration, which could happen if the live stream watch time is shorter.
@@
 Figure 2. Overall framework of the SL-MGAC algorithm. The SL (RL) MG-SD Module is short for the Multi-Group State Decomposition Module for supervised reward learning and critic learning.
-Overall, the optimization objective of aforementioned optimal live stream injection control problem is shown as follows:
+Overall, the concrete reward function $r(s_{t},a_{t})$ and constraint function $c(s_{t},a_{t})$ is shown below:
+(1) $r(s_{t},a_{t})$ $=y_{l}$
+$c(s_{t},a_{t})$ $=\frac{1}{B}y_{v}-y_{l}$
+where $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
+The optimization objective of aforementioned optimal live stream allocation control problem is shown as follows:
 (2) $\max_{\pi}J_{R}(\pi)$
 $s.t.\quad J_{C}(\pi)\leq\epsilon$
-$J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
-$J_{C}(\pi)=\frac{1}{B}y_{v}-y_{l}$
-where $\pi(a_{t}|s_{t})$ is the live stream injection policy to be optimized, $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
+where $J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, $J_{C}(\pi)=c(s_{t},a_{t}),\forall t\in[0,\infty)$, $\pi(s_{t},a_{t})$ is the live stream allocation policy to be optimized.
 A common method to solve the above CMDP is to transform the problem into a min-max optimization by introducing the Lagrange multiplier $\lambda$:
 (3) $(\pi^{*},\lambda^{*})=\arg\min_{\lambda\geq 0}\max_{\pi}J_{R}(\pi)-\lambda(J_{C}(\pi)-\epsilon)$
-Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function:
+Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function by substituting into $r(s_{t},a_{t})$ and $c(s_{t},a_{t})$ of Eq.
+1:
 (4) $\hat{r}(\lambda,s_{t},a_{t})$ $=r(s_{t},a_{t})-\lambda c(s_{t},a_{t})$
@@
 $=(1+\lambda)y_{l}-\frac{\lambda}{B}y_{v}$
-In practice, we solve the unconstrained MDP problem and the reward function is simplified as follows:
-(4) $r(s_{t},a_{t})=y_{l}-\frac{\lambda}{B}y_{v}$
-where $\lambda$ is a hyper-parameter to carefully control the final live stream injection ratio.
+In practice, we solve the unconstrained MDP problem with a simplified reward function as follows:
+(5) $\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(y_{l}-\frac{\lambda}{B}y_{v}\right)\right]$
+where $\lambda$ is a hyper-parameter to carefully control the final live stream allocation ratio.
 From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint.
+Note that we leave the further discussion of the rationality of the above reward function in the Appendix C.
 3. Proposed Framework
@@
 Otherwise, excessive live stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration.
-In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream injection policy for each individual user.
+In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream allocation policy for each individual user.
 Directly learning a policy from sparse and noisy user feedback is often infeasible, since it will inadvertently inject excessive live streams in the short video feed and highly affected the user engagement.
@@
 As shown in Eq.
-9, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin.
+10, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin.
 We then reconstruct the predicted live streaming and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively.
@@
 where $\psi$ is a random number, and $\epsilon$ is maximal exploration probability.
+Figure 3. System Architecture of the SL-MGAC algorithm.
 We implemented the SL-MGAC algorithm in our recommendation system, and the overall architecture of the system is shown in Fig.
@@
 Moreover, the offline trainer sends the latest model parameters to update the online deployed actor network in real-time.
-Figure 3. System Architecture of the SL-MGAC algorithm.
 4. Experiments
@@
 4.1. Dataset
-Due to the lack of publicly available datasets for recommendation decisions involving live streaming injections in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data.
+Due to the lack of publicly available datasets for recommendation decisions involving live streaming allocations in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data.
 In total, we have over 1,800,000 samples, which are divided into training and test sets with a 4:1 ratio.
@@
   • IQL (Kostrikov et al., 2021). An offline reinforcement learning algorithm that leverages expectile regression in Q-Learning.
-Note that in our optimal live stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
+Note that in our optimal live stream allocation control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
 4.2.2. Variations of our model
@@
 4.7. Online A/B Test Experiment
-We compare the improvements of different methods in terms of live stream daily active users (DAU), live stream watch time, video watch time, app usage duration and user retention relative to the baseline.
-The baseline uses the SAC framework for live stream injection.
+We compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline.
+The baseline uses the SAC framework for live stream allocation.
 The results of the online A / B test are shown in Table 3.
 Note that the Dummy method injects a live stream for each request.
-We can see that SL-MGAC achieves the largest improvement in live stream watch time while also increasing the app usage duration and user retention than other methods.
-It demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live streams in response to user requests.
-Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may negatively impact the long-term user experience for most users.
-Moreover, the dummy method can greatly improve live streaming DAU and watch time, but significantly affect app usage time and user retention.
+We can see that SL-MGAC achieves a significant improvement in live stream watch time while also increasing the app usage duration and user retention than other methods.
+This shows that SL-MGAC is more effective in maximizing long-term live stream rewards while avoiding the negative impact on app usage duration and user retention.
+Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may affect the long-term user experience for most users.
+Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage time and user retention.
 4.8. Online Model Stability
@@
 As shown in Fig.
-6, the live stream injection ratio fluctuates drastically throughout the day for both models.
+6, the live stream allocation ratio fluctuates drastically throughout the day for both models.
 However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude.
-Note that the live stream injection ratio refers to the proportion of requests with action $a=1$.
-Figure 6. The trend curve of online live stream injection ratio during a whole day.
-We calculate the amplitudes of the live stream injection ratio within sliding time windows of 20 minutes.
+Note that the live stream allocation ratio refers to the proportion of requests with action $a=1$.
+Figure 6. The trend curve of online live stream allocation ratio during a whole day.
+We calculate the amplitudes of the live stream allocation ratio within sliding time windows of 20 minutes.
 The amplitude density is shown in Fig.
@@
 In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, demonstrating the superior online model stability of SL-MGAC.
-5. Discussion
-The intrinsic instability and divergence of RL limits its application in real-world industrial recommendation systems.
-The proposed SL-MGAC method improves the robustness and stability of the RL model and achieves significant offline and online performance in the live stream and short video mixed recommendation scenario.
-However, SL-MGAC only learns from data from adjacent states $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1},r_{t+1})$, which limits its ability to learn information in much longer sequences, such as the daily-level user preference, rather than session-level user preference, since the capture of daily user preference information is very helpful for improving user engagement and user retention.
-Learning live stream policy from a long sequence of user behaviors would lead to our future work.
-Furthermore, we emphasize that more sophisticated methods can also be employed to divide users into multiple groups and provide prior information about user preference, such as traditional clustering methods, e.g.
-k-means, neural clustering of VQ-VAE (Van Den Oord et al., 2017), online clustering of SwAV (Caron et al., 2020) by the Sinkhorn-Knopp algorithm (Cuturi, 2013), etc.
-We find that the simple MG-SD module achieves significant performance and fast real-time response spped in practice; we will lead these complicated group partitioning methods in our further research.
 5. Related Work
@@
 6. Conclusion
-In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream injection policies in the challenging context of a live stream & short video mixed recommendation system.
-Unlike existing RL methods for recommendation systems, our approach focuses on enhancing the stability and robustness of the RL model.
+In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention in the challenging context of a live stream & short video mixed recommendation system.
 Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation.
-In practice, we aim to minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully deployed in large-scale industrial recommendation systems.
+Compared with existing RL methods for recommendation systems, our approach not only focuses on improving online performance, but also improves the stability and robustness of the RL model.
+In practice, we minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully implemented in large-scale industrial recommendation systems.
+Currently, we are trying to apply the SL-MGAC methods to other similar recommendation scenarios, including E-commerce and advertising.
+We also aim to continuously improve the performance of SL-MGAC in further research.
+Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
+7. Acknowledgment
+The authors acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.
 Appendix A Tables
@@
 Figure 8. The amplitude distribution of live stream injection ratio.
+Appendix C More Ablation Studies

### v5

@@
 Abstract.
-In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream into the video feed for each user request.
-To maximize long-term user engagement, it is crucial to determine an optimal live stream policy for accurate live stream allocation.
-The inappropriate live stream allocation policy can significantly affect the duration of the usage app and user retention, which ignores the long-term negative impact of live stream allocation.
+In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream to the video feed for each user request.
+The inappropriate policy which ignores the long-term negative impact of live stream allocation can significantly affect app usage duration and user retention.
+To maximize long-term user engagement, it is crucial to determine an optimal policy for accurate live stream allocation.
 Recently, reinforcement learning (RL) has been widely applied in recommendation systems to capture long-term user engagement.
-However, traditional RL algorithms often face divergence and instability problems, which restricts the application and deployment in the large-scale industrial recommendation systems, especially in the aforementioned challenging scenario.
+However, traditional RL algorithms often face divergence and instability problems, which restricts application and deployment in large-scale industrial recommendation systems, especially in the aforementioned challenging scenario.
 To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC).
-Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning.
+Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task supervised reward learning helps restrict bootstrapping error accumulation during critic learning.
 Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability.
@@
 Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing.
-Experimental results demonstrate that the proposed method not only outperforms baseline methods under the platform-level constraints but also exhibits enhanced stability in online recommendation scenarios.
+Experimental results demonstrate that the proposed method not only outperforms baseline methods under platform-level constraints, but also exhibits improved stability in online recommendation scenarios.
 Reinforcement Learning, Recommendation System, Variance Reduction
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2, August 3–7, 2025, Toronto, ON, Canada; Toronto, ON; Canadaisbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
 Figure 1. Structure of a short video & live stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live stream RS.
@@
 The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list.
-Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
+Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
 Currently, deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017).
@@
 In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
+Most importantly, the live-stream allocation agent is the final module of the live-stream RS, as shown in Fig 1.
+If the RL agent is very instable or collapse in the online environment, we may inject too many live-streams in the short-video feed, which will cause severe malfunctions and will highly affect the user retention.
+Hence, in the industrial RS, this RL application is more risky than existing RL methods (Cai et al., 2023; Zhang et al., 2022, 2024), which are only applied in the multi-rank score aggregation of the ranking stage.
 Furthermore, like advertisement, live stream is not welcomed by every user.
@@
 Additionally, we combine multi-task supervised reward learning with traditional critic learning to not only restrict temporal difference (TD) error accumulation during model training but also improve the accuracy of Q-value estimation.
-Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B testing to demonstrate its effectiveness and stability.
+Through experiments, we compare the SL-MGAC method with competitive baselines using offline policy evaluation (OPE) (Uehara et al., 2022) and online A/B tests to demonstrate its effectiveness and stability.
 The main contributions of this paper are as follows.
@@
   • We propose new variance reduction (Greensmith et al., 2004; Kumar et al., 2019; Anschel et al., 2017; Mao et al., 2018; Romoff et al., 2018) techniques, including multi-group state decomposition, distribution discretization for reward learning, reward normalization, and Q value normalization.
-  • We integrate multi-task supervised learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
+  • We integrate multi-task supervised reward learning with traditional critic learning to alleviate TD error accumulation and improve the accuracy of Q-value estimation.
   • We successfully deploy the SL-MGAC model in a challenging short video and live stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
@@
 Formally, we define live stream allocation CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
-  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, online viewer count).
+  • State space $\mathcal{S}$: This is the set of user interaction states $s$, which includes user static features (e.g., user ID, location, gender, country), user history features (e.g., live stream watch history, short video watch history) and item features (e.g., live stream ID, author ID, author gender, etc).
 We limit the length of user live stream and short video history lists by timestamp order, keeping only the top 50 items.
   • Action space $\mathcal{A}$: The action $a\in\mathcal{A}$ represents the decision of whether to inject a recommended live stream in response to a user’s request.
-We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream and $a=0$ means not injecting it.
+We define the action $a$ as a binary variable, where $a=1$ means injecting a live stream.
   • Transition Probability $\mathcal{P}$: The transition probability is denoted as $p(s_{t+1}|s_{t},a_{t})$, determined by the environment.
@@
 From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint.
-Note that we leave the further discussion of the rationality of the above reward function in the Appendix C.
+Note that we leave the further discussion of the above reward function in the Appendix C.
 3. Proposed Framework
@@
 3.3. Supervised Learning-enhanced Actor Critic
-To alleviate the impact of drastic changes in data distribution that could lead to the divergence or instability of RL models, we propose a supervised learning-enhanced actor critic framework to prevent critic networks from getting trapped in model collapse due to large cumulative errors in the critic learning process.
+To alleviate the impact of drastic changes in the data distribution that could lead to divergence or instability of RL models, we propose a supervised learning-enhanced actor critic framework to prevent critic networks from being trapped in model collapse due to large cumulative errors in the critic learning process.
 3.3.1. Layer Normalization
 A recent work on RL divergence and instability (Yue et al., 2024) shows that offline critic learning exhibits a self-excitation pattern.
-Specifically, Q-value iteration can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, which amplifies the TD-error and traps the critic learning process in a positive feedback loop until model collapse.
+Specifically, iteration of the Q value can inadvertently cause the target Q value $Q(s_{t+1},a^{*})$ to increase even more than the increment of $Q(s_{t},a_{t})$, amplifying the TD error and trapping the critic learning process in a positive feedback loop until the model collapse.
 Therefore, normalization techniques, such as Layer Normalization (Ba, 2016) can be utilized to alleviate divergence and instability problems.
@@
 $Q_{label}(s_{t+1})$ $=r(s_{t},a_{t})+\gamma\max_{a_{t+1}}\hat{Q}^{\prime}(s_{t+1},a_{t+1})$
-where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use Huber loss (Hastie et al., 2009) to optimize the above objective.
+where $\hat{Q}^{\prime}(s_{t+1},a_{t+1})=\min_{i=1,2}Q^{\prime}_{\phi_{i}}\left(s_{t+1},a_{t+1}\right)$. We use the Huber loss (Hastie et al., 2009) to optimize the above objective.
 In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of critic networks.
-Specifically, due to the instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term.
-Then the $r(s_{t},a_{t})$ term could not provide any information to guide the critic learning.
+Specifically, due to instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term.
+Then the term $r(s_{t},a_{t})$ could not provide any information to guide the learning of the critic.
 Therefore, we seamlessly introduce supervised learning (e.g.
@@
 Formally, we divide the live stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live stream (short video) watch time.
-Let $y$ denote the actual live stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$.
+Let $y$ denote the actual live stream or short video watch time, which falls in the time bin $[y_{st},y_{end}]$.
 The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$.
@@
 The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3.
-The discount factor $\gamma$ is set to 0.9 for all methods.
+The discount factor $\gamma$ is set to 0.9 for all methods and $\lambda$ in Eq 5 is 0.1.
 Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
@@
 IQL (Kostrikov et al., 2021) 432.03
+RLUR (Cai et al., 2023) 443.01
 SL-MGAC-sep 435.19
@@
 Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may affect the long-term user experience for most users.
-Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage time and user retention.
+Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage duration and user retention.
 4.8. Online Model Stability
@@
 The work in (Miao et al., 2021) explores the use of off-policy RL for personalized search ranking in multiple sessions.
-CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving the user experience.
+CrossDQN (Liao et al., 2022) introduces an RL-based approach to ad allocation in a feed, aiming to maximize revenue while improving the user experience.
 Furthermore, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
@@
 6. Conclusion
-In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention in the challenging context of a live stream & short video mixed recommendation system.
-Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation.
+In the challenging context of a live stream & short video mixed recommendation system, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention.
+Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task supervised reward learning with traditional critic learning to constrain Q-value estimation.
 Compared with existing RL methods for recommendation systems, our approach not only focuses on improving online performance, but also improves the stability and robustness of the RL model.
@@
 Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
-7. Acknowledgment
-The authors acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.
 Appendix A Tables
@@
 Appendix C More Ablation Studies
+We conduct further online AB experiments to demonstrate the performance of the SL-MGAC policy at different values of $\lambda$ in Eq 5, as shown in Table 5.
+It indicates that when $\lambda>0$, the live stream allocation policy of SL-MAGC behaves more conservatively to balance live stream watch time and video watch time, and finally optimize user’s long-term engagement and retention.
+However, when $\lambda<0$, the SL-MAGC behaves more greedily to allocate live streams to the video feed.
+Specifically, we find that the SL-MAGC allocates more live streams to the requests that occur earlier in a session due to the position bias on the reward, which would interrupt the user interest throughout the session and hence affect user retention.
+$\lambda$ live stream DAU live stream Watch Time Video Watch Time App Usage Duration User Retention
+0.2 -1.245% -3.567% +0.215% +0.123% +0.106%
+0.1 +2.616% +7.431% +0.197% +0.121% +0.086%
+-0.1 +4.123% +6.564% -0.248% -0.362% -0.112%
+-0.2 +5.876% +7.215% -0.296% -0.445% -0.159%
+Table 5. Online AB Test Performance of SL-MGAC under different values of $\lambda$.

### v6

@@
 Reinforcement Learning, Recommendation System, Variance Reduction
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2, August 3–7, 2025, Toronto, ON, Canada; Toronto, ON; Canadaisbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737264isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning
 Figure 1. Structure of a short video & live stream mixed recommendation system(RS). The decision making of SL-MGAC takes place in the final stage of live stream RS.
@@
 Most importantly, the live-stream allocation agent is the final module of the live-stream RS, as shown in Fig 1.
-If the RL agent is very instable or collapse in the online environment, we may inject too many live-streams in the short-video feed, which will cause severe malfunctions and will highly affect the user retention.
+If the RL agent becomes highly unstable or collapses in the online environment, it may lead to an excessive injection of live-stream content into the short-video feed.
+In turn, this can negatively impact the exposure of advertisements and E-commerce videos, potentially causing severe system malfunctions and significant platform losses.
 Hence, in the industrial RS, this RL application is more risky than existing RL methods (Cai et al., 2023; Zhang et al., 2022, 2024), which are only applied in the multi-rank score aggregation of the ranking stage.
@@
 Therefore, we cannot merely focus on the user feedback for a single request, but rather should optimize the long-term user experience.
-To address the aforementioned issues, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC.
+To address the aforementioned problems, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm, SL-MGAC.
 Given the significant variation in users’ interest in live stream content, which introduces high variance at both the feature and feedback levels, we incorporate separate Multi-Group State Decomposition (MG-SD) modules within both the actor and critic networks.
@@
   • We successfully deploy the SL-MGAC model in a challenging short video and live stream mixed recommendation scenario for Kwai, a short video app with over 100 million users.
+Figure 2. Overall framework of the SL-MGAC algorithm. The SL (RL) MG-SD Module is short for the Multi-Group State Decomposition Module for supervised reward learning and critic learning.
 2. Problem Formulation
@@
 By meeting this constraint, we ensure that live streams boost users’ long-term app engagement instead of reducing their time spent on it.
-Figure 2. Overall framework of the SL-MGAC algorithm. The SL (RL) MG-SD Module is short for the Multi-Group State Decomposition Module for supervised reward learning and critic learning.
 Overall, the concrete reward function $r(s_{t},a_{t})$ and constraint function $c(s_{t},a_{t})$ is shown below:
@@
 We find that SL-MGAC without the MG-SD module tends to be more unstable during training, with a larger variance in its Q-values.
+Furthermore, the MG-SD module is flexible and transferable, allowing the user-group partitioning strategy to be easily adapted to various scenarios in advertising and E-commerce.
 3.3. Supervised Learning-enhanced Actor Critic
@@
 Moreover, the offline trainer sends the latest model parameters to update the online deployed actor network in real-time.
+Most importantly, we focus on reducing computational costs for successful deployment in two key aspects.
+First, for real-time training, we avoid using user-item cross features, which are commonly employed in industrial ranking models, thereby reducing the sparse embedding parameters.
+Additionally, the network parameters of the Actor part in SL-MGAC are much smaller than those of the Critic, with the complex Critic acting as a teacher to guide the Actor’s learning, as explained in Section 3.3.3.
+As a result, the total parameters of SL-MGAC are under 100 million in our recommendation system, significantly smaller than the parameter scale of typical ranking models, which usually exceed 1 billion parameters.
+For fast online inference, the smaller Actor network enables fast inference, processing each request in under 20ms, even with over 10,000 requests per second during peak times.
+Furthermore, since the complex Critic is not used during online inference, this further reduces computational costs, contributing to faster response time in the online environment.
 4. Experiments
@@
   • IQL (Kostrikov et al., 2021). An offline reinforcement learning algorithm that leverages expectile regression in Q-Learning.
+  • RLUR (Cai et al., 2023). An RL model that aims to optimize the long-term user retention.
 Note that in our optimal live stream allocation control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
@@
 Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
+And the code of SL-MGAC method is available at https://github.com/frankg1/SL-MGAC-torch.
 4.4. Offline Policy Evaluation
@@
 4.
+Note that the shaded area in yellow corresponds to the Q value std range of the SL-MGAC model, while the shaded area in purple corresponds to the Q value std range of the SL-MGAC model(w/o MG).
 We observe that the Q-value curve of SL-MGAC is more stable than that of SL-MGAC (w/o MG), and the Q-value variance of SL-MGAC is much smaller.
 This demonstrates the effectiveness and variance reduction effect of the MG-SD module.
-Figure 4. The Q value curves between SL-MGAC and SL-MGAC (w/o MG) over 10 rounds of training. The lines correspond to the means of Q-value and the shaded areas correspond to the standard deviations.
+Figure 4.
+The Q value curves between SL-MGAC and SL-MGAC (w/o MG) over 10 rounds of training.
+The lines correspond to the means of Q-value and the shaded areas correspond to the standard deviations (std).
 4.6. Parameter Sensitivity
@@
 4.7. Online A/B Test Experiment
-We compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline.
+The online A/B Test is conducted over a 5-day period in 2024 September and we randomly choose 20% users (over 20 million) in Kwai app as the experimental group.
+Then we compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline.
 The baseline uses the SAC framework for live stream allocation.
@@
 We evaluate the online model stability between the SAC-based baseline and SL-MGAC.
+The live-stream injection ratio is chosen as the evaluation metric, calculated by dividing the number of live-stream injection requests by the total number of requests within a given timestamp.
 As shown in Fig.
-6, the live stream allocation ratio fluctuates drastically throughout the day for both models.
+6, the live-stream allocation ratio for both models fluctuates significantly throughout the day, from high-traffic hours (such as 6:00–11:00) to low-traffic hours (14:00–18:00).
+This variation is due to changes in the overall user scale during these periods, which in turn affect the data distribution.
 However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude.
@@
 Their supervised learning task focuses on predicting the next item recommendation probability, whereas our approach introduces supervised learning to restrict critic learning.
+However, the aforementioned RL methods may encounter instability issues and could fail when faced with highly noisy or drastically changing data distributions.
+The proposed SL-MGAC method is more robust and has been successfully deployed in the ”risky” final stage of a real-world industrial RS.
+We believe this work provides valuable insights not only into the model framework design for RL applications in industrial RS, but also offers practical experience in deploying robust and successful RL systems for other industrial companies.
 5.2. RL divergence and instability
@@
 Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
+7. Acknowledgment
+The authors acknowledge Wei Bai, Xiaoshuang Chen and anonymous reviewers for proposing detailed modification advice to help us improve the quality of this manuscript.
 Appendix A Tables